3. The phenotypic variance for litter size is 5.85 pigs2 and the heritability of litter size is 0.15. A sow produces a litter of 13 piglets in a contemporary group that averages 10 piglets per litter per sow.

A) What is this sow’s breeding value?

B) What are the 68% and 95% confidence ranges for this sow’s breeding value? Interpret these confidence ranges.
Answer #1

A) litter of piglets in contemporary group - average piglets per sow = 13 - 10 = 3

heritability of litter size , h^2 = (0.15)^2 = 0.0225

sow breeding value = 3* 0.0225 = 0.0675.

B) With 68% and  95% confidence interval are analysed  to measure of return to a recurring period of sexual receptivity and fertility after first insemination, and the controlling for the selection of the pigs by incidence of return to sexual receptivity and fertility recurring period before entry into the study.